OpenStudy (anonymous):

Simplify: i^41 1) i 2) -1 3) -i 4) 1

OpenStudy (anonymous):

1) i

OpenStudy (anonymous):

Simplify: i^72 1) i 2) -1 3) -i 4) 1

OpenStudy (anonymous):

4) 1

myininaya (myininaya):

i i^2=-1 i^3=i*i^2=i*(-1)=-i i^4=i^2*i^2=1 i^5=i^4*i=1*i=i i^6=i^5*i=i*i=i^2=-1 this is gonna keep repeating it self i,-1,-i,1,i,-1,-i,1, and so on... the trick is to take the 72 and figure out the remainder when 72 is divided by 4

OpenStudy (anonymous):

So would this answer i or 1?

myininaya (myininaya):

18 ______ 4| 72 -4 ---- 32 -32 ---- 0 the remainder is 0 so you have \[i^{72}=i^{4(18)+0}=(i^4)^{18}=(1)^{18}=1\] ---------------------- if you have \[i^a\] \[i^a=i^{4(k)+j}\] if j=0, then answer is 1 if j=1, then answer is i if j=2, then answer is -1 if j=3, then answer is -i

myininaya (myininaya):

assuming a=4k+j of course

OpenStudy (anonymous):

So the answer is #1?

myininaya (myininaya):

if you read what i have then you will find the answer i did it for you and so did euler i was just showing work